# LUMA: A Benchmark Dataset for Learning from Uncertain and Multimodal Data

### v2

@@ Appendix C Details about Baseline architectures and Datasets
Parameter Clean Dataset Reduced Diversity Increased Label Noise Increased Sample Noise
Compactness 0 20 0 0
Sample Noise False False False True
Label Switch Prob. 0 0 0.3 0
Noisy Data Ratio 0.0 0.0 0.0 1
Audio Noise SNR None None None 3-5
Image Noise None None None
Gaussian Noise: 4,
Shot Noise: 4,
Impulse Noise: 4,
Defocus Blur: 4,
Frosted Glass Blur: 4,
… (additional noises
summarized)
Text Noise None None None
KeyboardNoise: 1-5 char,
3-8 word,
BackTranslation,
Spelling: 0.4,
OCR: 0.5 word
Table 3: The parameters for generating the different versions of the dataset for experiments with the baseline models.
C.1 Dataset Generation
We compiled one clean version and three noisy versions of the dataset, each with a specific type of noise: reduced diversity, increased label noise, and increased sample noise. The parameters for compiling the different versions can be found in Table 3. These parameters were selected to introduce sufficient noise, ensuring observable changes in the uncertainty metrics. Please refer to our GitHub directory for the full configuration files.
C.2 Baseline models
In this section, we will describe the baseline models employed for evaluating the dataset. These baselines were selected to instantiate unimodal and multimodal architectures, which can be trained on the dataset and are not intended to serve as a comprehensive benchmark, nor did we endeavor to achieve the best possible performance.
@@ C.2 Baseline models
For the image modality, we used a simple convolutional neural network, as depicted in Figure 11. For the audio modality, we extracted 128x128 mel-spectrograms from padded audio samples, and used a convolutional network for classification, as depicted in Figure 1. For the text modality, we extracted the BERT (Devlin et al., 2019a) embeddings for each token, and averaged them out, so that we have one embedding per text passage. Then, we passed the embedding through a simple feed-forward neural network (Figure 13) to get the predictions. As depicted in Figures 11, 12 and 13, each model includes two output heads: one for the prediction and the other for aleatoric uncertainty, following the methodology outlined by Valdenegro-Toro & Mori (2022).
Figure 11: The classification network for the image modality.
Figure 12: The classification network for the audio modality.
Then, to combine the aforementioned unimodal networks into a multimodal architecture, we adopted the late fusion approach. In the Monte Carlo Dropout and Deep Ensemble methods, we obtained the multimodal prediction by averaging the logits from the final layers of the classifiers. For the RCML, we modified the output of the last layer in each network to produce evidence, as described in (Xu et al., 2024), and followed their methodology for combining the evidence.
In Figures 11, 12 and 13 the Dropout layer is kept turned on during the inference stage, for Monte-Carlo Dropout method, and turned off for Deep Ensemble and RCML methods. The dropout probability for all models is 0.3. The number of networks in deep ensembles is 10.
Figure 13: The classification network for the text modality.
@@
We trained the networks for up to 300 epochs, with early stopping applied after 10 epochs of no improvement in validation loss. The learning rate was set to 0.001 and was reduced by a factor of 0.33 if there was no improvement in validation loss after 5 epochs.
The experiments were conducted on an internal cluster equipped with NVIDIA GeForce RTX 2080 GPUs, totaling approximately 22 GPU hours.
C.3 Classification accuracies for baseline models
In Table 4 we present the classification accuracy measures for the clean dataset and variations in accuracy under different types of noise for the three baseline models (Monte Carlo Dropout, Deep Ensemble, and RCML) across each modality. Each version of the noisy dataset incorporates a single type of noise: Reduced Diversity (K=20), Label Noise (30%), and Sample Noise, with severity varying by noise type and modality (detailed in supplementary material). The RCML model, a multimodal approach, is evaluated exclusively in the multimodal setting. We observe that accuracy always decreases with increasing the label noise, but reducing diversity and increasing sample noise may not always decrease accuracy in the image modality.
@@ C.3 Classification accuracies for baseline models
Table 4: Classification accuracies for the clean dataset and variations in accuracy under different types of noise.

### v3

@@ 4. Baseline models with Uncertainty Quantification
Method Clean Dataset Reduced Diversity Increased Label Noise Increased Sample Noise
Ale. Epi. Ale. Epi. Ale. Epi. Ale. Epi.
MCD Image 1.00 1.03 -15.73% -11.66% +59.20% +54.51% +4.44% +2.18%
MCD Audio 0.52 0.70 -5.54% +2.16% +96.63% +54.49% +23.12% +14.40%
MCD Text 0.37 1.01 -3.91% -2.62% +93.59% +2.41% +64.96% -2.03%
MCD Multi. 0.26 0.78 -8.52% -1.21% +122.44% +11.60% +59.14% +9.89%
DE Image 1.45 1.40 -37.49% -8.54% -7.43% +0.24% -18.46% -3.22%
DE Audio 0.56 0.99 -27.39% -3.34% +156.40% +50.43% +70.26% +34.41%
DE Text 0.42 1.01 +5.02% -6.15% +81.26% -0.51% +62.24% -7.11%
DE Multi. 0.31 0.82 -22.80% -3.40% +115.15% +20.62% +45.97% +5.54%
Table 2. OOD Detection AUC Values for Different Methods
Method MCD DE RCML Multi.
Image Audio Text Multi. Image Audio Text Multi.
AUC 0.54 0.47 0.53 0.50 0.54 0.49 0.54 0.50 0.91
We develop baseline models with three different uncertainty quantification algorithms, to serve as a starting point for other research and benchmarking initiatives. For the sake of simplicity, we choose late fusion approaches, where we have classification networks for each modality, and then fuse their decisions by simply averaging the output logits. These baselines were selected to instantiate unimodal and multimodal architectures, which can be trained on the dataset and are not intended to serve as a comprehensive benchmark, nor did we endeavor to achieve the best possible performance. The architectures of the baseline models are depicted in Figure 9.
Figure 9. Network architectures used for image, audio and text modalities.
[…]
Figure 10: Fully described in text.
@@
For the image modality, we used a simple convolutional neural network. For the audio modality, we extracted 128x128 mel-spectrograms from padded audio samples, and used a convolutional network for classification. For the text modality, we extracted the BERT (Devlin et al., 2019b) embeddings for each token, and averaged them out, so that we have one embedding per text passage. Then, we passed the embedding through a simple feed-forward neural network to get the predictions. As depicted in Figure 9, each model includes two output heads: one for the prediction and the other for aleatoric uncertainty, following the methodology outlined by Valdenegro-Toro and Mori (2022). Then, to combine the aforementioned unimodal networks into a multimodal architecture, we adopted the late fusion approach. In the Monte Carlo Dropout and Deep Ensemble methods, we obtained the multimodal prediction by averaging the logits from the final layers of the classifiers. For the Reliable Conflictive Multi-View Learning (RCML) (Xu et al., 2024), we modified the output of the last layer in each network to produce evidence, as described in (Xu et al., 2024), and followed their methodology for combining the evidence.
@@
The dropout probability is 0.3, with the deep ensemble comprising 10 networks. Networks were trained for up to 300 epochs, with early stopping after 10 epochs of no validation loss improvement.
@@
Table 3. Classification accuracies for the clean dataset and variations in accuracy under different types of noise.
@@
4.1. Uncertainty Metrics
For uncertainty quantification, we implemented 3 approaches: Monte Carlo Dropout (MCD) (Gal and Ghahramani, 2016), Deep Ensemble (DE) (Lakshminarayanan et al., 2017), Reliable Conflictive Multi-View Learning (RCML) (Xu et al., 2024). In Monte Carlo Dropout and Deep Ensembles, we use the aleatoric entropy and the epistemic entropy as uncertainty measures, and follow Valdenegro-Toro and Mori (2022) for disentangling the aleatoric and epistemic uncertainties. $H_{\text{Ale }}(y\mid\mathbf{x})=\operatorname{entropy}\left(p_{\text{Ale }}(y\mid\mathbf{x})\right)$ and $H_{\mathrm{Epi}}(y\mid\mathbf{x})=\operatorname{entropy}\left(p_{Epi}(y\mid\mathbf{x})\right)$, where $p_{Epi}$ and $p_{Ale}$ are the probabilities obtained according to (Valdenegro-Toro and Mori, 2022). For RCML (Xu et al., 2024), we measure the aleatoric uncertainty with the expected entropy:
(2) $\mathbb{E}_{p(\boldsymbol{\pi}\mid\mathbf{x},\hat{\boldsymbol{\theta}})}[H[P(y\mid\boldsymbol{\pi})]]=-\sum_{k=1}^{K}\frac{\alpha_{k}}{\alpha_{0}}\left(\psi\left(\alpha_{k}+1\right)-\psi\left(\alpha_{0}+1\right)\right),$
where $\alpha_{k}$ is the $k$-th concentration parameter of the Dirichlet distribution, and $\alpha_{0}$ is the sum of all concentration parameters. $\psi$ is the digamma function. As a measure for epistemic uncertainty, we take $\frac{N}{\alpha_{0}}$, where $N$ is the number of classes. We evaluate the measures of accuracy and uncertainty of the models on the clean dataset, the dataset with reduced diversity, the dataset with increased sample noise, and the dataset with increased label noise [^12] For noise generation parameters, please refer to our GitHub page.
4.2. Results
The results are summarized in Table 1. Since uncertainty is quantified differently in RCML than in MCD and DE, their values cannot be directly compared, so the table reports the relative changes in both types of uncertainty with respect to the clean dataset.
We evaluate AUC score for OOD detection based on the epistemic uncertainty. The results are summarized in Table 2. We can see that Monte Carlo Dropout and Deep Ensembles fail to provide epistemic uncertainty values suitable for OOD detection in LUMA dataset, with a poor performance of approximately 0.5 AUC value. On the other hand, the RCML achieves an outstanding AUC score of 0.91, indicating that the epistemic uncertainty values quantified with this method can be effectively used for OOD detection.
To further evaluate the qualities of the uncertainties of the different models, we estimate the epistemic and aleatoric uncertainties under different amounts of label noise. Ideally, we expect a good uncertainty quantification algorithm to provide higher uncertainty values for more noisy data. As we can see from Figure 10, only RCML consistently raises the uncertainty estimates under increased label noise, which again shows the higher quality of its uncertainty estimates over the other baselines.
In Table 3, we present the classification accuracy measures for the clean dataset and variations in accuracy under different types of noise for the three baseline models across each modality. We observe that accuracy always decreases with increasing the label noise, but reducing diversity and increasing sample noise may not always decrease accuracy in the image modality.
Figure 10. Changes in uncertainty estimations under different proportions of label noise (in percentages). The RCML approach consistently increases both aleatoric and epistemic uncertainties with increased label noise. In contrast, the MCD and DE models sometimes fail to increase the corresponding uncertainty estimations in this experiment.
[…]
Figure 10: Fully described in text and caption.
5. Conclusion
In this paper, we propose LUMA, a multimodal dataset for learning from uncertain and multimodal data. The dataset spans image, audio, and text modalities, and is accompanied by a Python package that allows users to generate customized versions with varying levels and types of noise and uncertainty. To support future research, we also provide a suite of baseline implementations for performance comparison. The dataset can be easily extended with additional modalities and augmented with more data samples. In the future, we plan to include dependent modalities, enabling more comprehensive studies of uncertainty quantification as well as improved applicability in information retrieval contexts. The open-source nature of the data compilation pipeline and code for uncertainty and noise generation facilitates the integration of new contributions from the community to promote multimodal uncertainty studies and benchmarking initiatives.
Acknowledgements.
We thank our colleagues and friends for contributing to the audio validation process, and to the HumanSignal team for providing an academic license for Label Studio.